OpenStudy (anonymous):

Write a quadratic function in intercept form whose graph has given x- intercepts and passes through the given point. x-intercepts: -2,2 Point: (-4,8)

hero (hero):

Intercept form of quadratic function is: f(x) = a(x - p)(x - q) where p and q are the intercepts: So we have: f(x) = a(x - 2)(x + 2) Now we just have to find the value of a. We can simply input the point (-4,8) to find a: 8 = a(-4 - 2)(-4 + 2) 8 = a(-6)(-2) 8 = 12a 8/12 = a 2/3 = a Thus the quadratic function given point (-4,8) and intercepts 2 and -2 is f(x) = 2/3(x - 2)(x + 2)